Fermi National Laboratory
 

Operation Summary for January 18, 2010 - January 20, 2010

 

Monday Jan. 18

The day shift began with Operations monitoring store 7512 (126.63E30) and the stash (95.2E10), with the Antiproton Source (Pbar) stacking (23.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and MTest T977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, at 10:38 AM, Operations reported a MI vacuum burst.

During the evening shift, Operations reported a 750 KeV line quadrupole power supply inhibit that halted beam from 8:45 to 10:15 PM due to overheating. NuMI personnel came in and investigated a MI-65 neutralization water system problem. Operations resumed sending beam to NuMI at 11 PM. Operations terminated store 7512 at 11 PM. Operators then made TeV orbit corrections.

Tuesday Jan. 19

The midnight shift began with Operations monitoring the stash (415.7E10), with Pbar stacking (31.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, a Linac Klystron RF station (KRF4) tripped off at 12:55 AM due to a LCW regulation loop problem. Operators restored Linac beam by 3:20 AM. Operations put the TeV into HEP shot setup at 3:35 AM and established store 7516 at 5:30 AM with an average initial luminosity (AIL) of 277.33E30. Operations established beam to MTest at 6 AM.

During the day shift, a Linac quadrupole power supply (L:QPS505) tripped off at 12:40 PM and required a power supply replacement.

During the evening shift, Linac personnel completed the quadrupole power supply replacement at 4:30 PM. Operations disabled the MTest beam at 6 PM. Operations terminated store 7516 at 7 PM and then prepared D0 for an access. After D0 came out of access, Operations put the TeV through dry and wet squeezes.

Wednesday Jan. 20

The midnight shift began with Operations monitoring the stash (403.1E10), with Pbar stacking (28.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TeV into shot setup at 12:07 AM and established store 7520 at 1:45 AM with an AIL of 269.01E30. Operations established beam to MTest at 4:30 AM. Operations reported a vacuum burst in the MI RF section at 7:20 AM; the vacuum had recovered by 7:45 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Linac personnel to flush their H- Source water resistor this afternoon.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/03/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory